H_{3}^{+} IN DIFFUSE INTERSTELLAR CLOUDS: A TRACER FOR THE COSMIC-RAY IONIZATION RATE

\underline{NICHOLAS INDRIOLO}, BENJAMIN J. McCALL, Department of Astronomy and Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720; TAKESHI OKA, Department of Astronomy & Astrophysics and Department of Chemistry, University of Chicago, Chicago, IL 60637.

In the past several years H_{3}^{+} has been detected^{a} in diffuse interstellar clouds where it had been predicted to exist in abundances below observable limits. These detections led to several questions about H_{3}^{+} chemistry and the diffuse cloud environment. To answer these questions, further studies of H_{3}^{+} were performed to constrain some uncertain parameters such as the electron recombination rate constant^{b}. With new data and some reasonable assumptions it became possible to infer the cosmic-ray ionization rate necessary to produce column densities of H_{3}^{+} observed by the CGS4 spectrometer on the United Kingdom Infrared Telescope (UKIRT). However, the value calculated along the sightline to ζ Per was over an order of magnitude higher than the value found in previous studies using OH^{c}^{,}^{d}.

Recently, we finished analyzing a survey of H_{3}^{+} in twenty diffuse cloud sightlines. In total, eight of these sightlines yielded positive detections of H_{3}^{+} with another one or two hinting at possible absorption. Using the observed column densities, we calculated the primary cosmic-ray ionization rate and found an average value of ζ_{p} ≈ 2 × 10^{−16} sec^{−1}. This is about an order of magnitude larger than what has long been considered the canonical ionization rate. Taking this higher ionization rate into account, some newer models^{e} have been proposed to explain both the H_{3}^{+} column densities, as well as the abundances of other species. The observation of H_{3}^{+} along many diffuse cloud sightlines, coupled with the lack of a clear correlation between position and ionization rate, suggests that an enhanced cosmic-ray ionization rate may be a general property of the diffuse interstellar medium.

aB. J. McCall, et al., ApJ, 567, 391 (2002)

bB. J. McCall, et al., Nature, 422, 500 (2003)

cJ. H. Black, T. W. Hartquist, & A. Dalgarno, ApJ, 224, 448 (1978)

dS. R. Federman, J. Weber, & D. L. Lambert, ApJ, 463, 181 (1996)

eH. Liszt, A&A, 398, 621 (2003)